Find the EMF and resistance of the battery if the current in the external circuit is 300 mA, the
net power of the circuit is 1.8 W, and the current in the external circuit is 100 mA, it is 0.1 W.
1
Expert's answer
2020-11-23T05:30:59-0500

The power is

"P=I^2(R+r),"

where I - current in the circuit;

R - resistance of external circuit;

r - internal resistance of battery.

For two cases above:

"1.8=(0.3)^2\cdot (R+r),\\
0.1=(0.1)^2 \cdot (R+r)."

From first equation

"R+r=20"

From second equation

"R+r=10."


I believe that the condition of the problem is incorrect